1. A capacitor consists of two square metal plates, each measuring 5.0 x 10-2 m on a side. In between the plates is a sheet of mica measuring 1.0 x 10-4 m thick. (a) What is the capacitance of this capacitor? (b) If the charge in one plate is 2.0 x 10-8 C, what is the potential difference between the plates?
1
Expert's answer
2022-04-29T12:30:18-0400

(a)

"C=\frac{\kappa \epsilon_0A}{d}\\
=\frac{5.4*8.85*10^{-12}*5.0*10^{-2}}{1.0*10^{-4}}=2.4*10^{-8}\:\rm F"

(b)

"V=\frac{q}{C}=\frac{2.0*10^{-8}}{2.4*10^{-8}}=0.83\:\rm V"